# In-context KV-Cache Eviction for LLMs via Attention-Gate

### v2

@@ Abstract
The KV-Cache technique has become the standard for the inference of large language models (LLMs). It caches states of self-attention to avoid recomputation. Yet, it is widely criticized that KV-Cache can become a bottleneck of the LLM inference system, especially when confronted with ultra-large models and long-context queries. A natural remedy is to discard the KV-Cache for less important tokens, with StreamingLLM (Xiao et al., 2024) as an example, but the used static eviction strategies cannot flexibly adapt to varying contexts. Remedies like H2O (Zhang et al., 2024) leverage accumulative attention scores to perform dynamic eviction but suffer from the attention bias issue in capturing contextual information. This paper bridges this gap by devising a parameterized KV-Cache eviction mechanism, dubbed as Attention-Gate, which accepts the whole context as input and yields eviction flags for each token to realize in-context eviction. The subsequent self-attention module proceeds according to the flags and only the KV states for the remaining tokens need to be cached. The Attention-Gates can vary among different heads and layers and be trivially plugged into pre-trained LLMs, tuned by cost-effective continual pre-training or supervised fine-tuning objectives to acquire what to discard. The computational and memory overhead introduced by Attention-Gates is minimal. Our method is validated across multiple tasks, demonstrating both efficiency and adaptability. After a highly efficient continual pre-training, it achieves higher average accuracy and evicts more tokens compared to traditional training-free methods. In supervised fine-tuning, it not only evicts many tokens but also outperforms LoRA-finetuned LLMs on some datasets, such as RTE, where it improves accuracy by 13.9% while evicting 62.8% of tokens, showing that effective eviction of redundant tokens can even enhance performance.
@@ 1 Introduction
Large language models (LLMs) (Dubey et al., 2024; Team et al., 2024; Chiang et al., 2023) have achieved remarkable success across a wide range of natural language processing tasks. A key technique that has enabled efficient LLM inference is KV-Cache, which stores transient attention keys and values to avoid recomputation. However, as the size of LLMs continues to increase and the demand for handling long-context queries grows, the KV-Cache has emerged as a significant bottleneck. Storing attention states for numerous tokens can lead to considerable memory overhead and data transfer among the memory hierarchy results in substantially increased inference time.
@@
To overcome these challenges, we introduce a parameterized KV-Cache eviction mechanism named the Attention-Gate (AG), to perform reliable in-context eviction. AG is positioned before a self-attention layer within the model. It takes a sequence of token features as input and generates eviction flags for the tokens, indicating whether a token should be excluded from subsequent self-attention computations. Tokens that are evicted do not require their KV states to be cached. AGs can be seamlessly integrated into pre-trained LLMs and tuned by minimizing the language modeling loss. Ideally, AGs can automatically learn to discern the most relevant tokens for the current context without manual intervention. In practice, we can implement the AG as a self-attention layer with much fewer heads than the original model (e.g., 4 v.s. 32). This way, the parallel computational capabilities of the hardware can be harnessed to minimize the extra overhead introduced by AGs.
AG is empirically shown to enjoy high training efficiency, e.g., only four NVIDIA 4090 GPUs and a dataset of 5,000 samples are required for continual pre-training when applying AGs to LLaMA2-7B (Touvron et al., 2023). This alleviates concerns about the computational overhead related to trainable eviction strategies (Zhang et al., 2024; Chen et al., 2024) and amplifies the performance superiority of our approach over existing training-free approaches. As illustrated in Figure 1, AG generates different eviction strategies across different layers and attention-heads for different tokens, demonstrating its adaptability to the diverse requirements of each component in the model.
To validate the effectiveness of our method, we conduct extensive experiments across multiple benchmarks. After efficient continual pre-training (CPT), our approach outperforms traditional training-free eviction strategies, such as StreamingLLM and H2O, in both accuracy and token eviction rates. In supervised fine-tuning (SFT), our method not only evicts a significant number of redundant tokens but also maintains or surpasses the performance of LoRA-finetuned LLMs. For example, on the RTE dataset, our approach improves accuracy by 13.9% while evicting 62.8% of tokens, demonstrating that selective token eviction can enhance performance. In summary, the Attention-Gate mechanism provides a scalable and efficient solution for KV-Cache management, addressing the limitations of traditional training-free methods.
@@ 2 Related Work
Computation and memory challenges. The computational cost of attention is driven by its quadratic complexity, while the memory burden stems from the need to store the KV-Cache during inference to avoid recomputation. Although several works have focused on reducing the computational load of attention mechanisms, many fail to address the memory constraints that remain a critical issue. Techniques like Reformer (Kitaev et al., 2020) and FlashAttention (Dao et al., 2022; Dao, 2023) mitigate memory usage for long sequences but still require substantial cache storage. On the other hand, approaches such as MQA (Shazeer, 2019) and GQA (Ainslie et al., 2023) prioritize memory compression without tackling computational complexity. These solutions tend to focus on either memory or computational efficiency, rarely addressing both simultaneously.
KV-Cache eviction strategies. To address both memory and computational challenges, KV-Cache eviction has emerged as an effective strategy. Existing approaches can be categorized into static and adaptive methods based on accumulative attention scores.
@@
Strategies using accumulative attention scores offer more flexibility by dynamically identifying important tokens. For instance, SpAtten (Wang et al., 2021) employs Accumulative Attention Scores (A2S), which sum the softmax outputs for each token to measure its importance. This approach allows selective token pruning in subsequent layers, effectively reducing computational complexity without the need for retraining. H2O (Zhang et al., 2024) extends this concept to decoder-based models, using local A2S statistics for adaptive eviction in autoregressive generation. However, H2O suffers from the attention bias issue (Oren et al., 2024), particularly in long-context inputs. Several follow-up works have aimed to address this limitation. NACL (Chen et al., 2024) introduces random eviction to mitigate attention bias, while A2SF (Jo & Shin, 2024) incorporates a Forgetting Factor. However, none of these approaches fully resolve the underlying problem.
@@ 3.1 Preliminary
Multi-Head Attention (MHA) (Vaswani et al., 2017) is a core component of the Transformer architecture, widely used in most LLMs. MHA enables the model to capture dependencies across different tokens in a sequence by computing self-attention across multiple attention-heads. Each head attends to different parts of the input sequence independently, allowing the model to capture various aspects of the relationships between tokens. The outputs of these attention heads are then concatenated and projected through a final output matrix.
KV-Cache is employed to store the key and value representations of tokens from previous time steps during inference. This prevents redundant recomputation of these representations for every new token generated, significantly speeding up auto-regressive generation. The inference process in auto-regressive transformers, such as Llama (Touvron et al., 2023), can be divided into two stages: prefilling and decoding. 1. In the prefilling stage, the model processes the entire input sequence, generating key-value pairs for each attention-head in each layer. These key-value pairs are stored in the KV-Cache for future reuse. 2. In the decoding stage, when generating the next token, the model only computes the key and value for that new token and appends them to the cached key-value pairs. This reduces the computational load, as the previously cached key-value pairs do not need to be recomputed. The process continues token by token until the sequence generation is complete.
KV-Cache plays a critical role in improving the efficiency of LLM inference, especially in scenarios where long sequences are processed. However, the size of the KV-Cache grows with the input sequence length, leading to substantial memory overhead. Efficiently managing this cache while maintaining model performance has become a key challenge in scaling LLMs to longer contexts.
@@ 3.2 Limitations of Traditional Eviction Strategies
Lack of flexibility. Flexibility in KV-Cache eviction strategies spans across various dimensions, including token-specific, attention-head-specific, layer-specific, task-specific, and model-specific adaptability. Static eviction policies offer no such flexibility, requiring manual intervention to adjust them. While methods based on accumulative attention scores (e.g., H2O (Zhang et al., 2024)) have improved upon static strategies by introducing token-level and head-level adaptability, they are still limited. Specifically, in H2O, the eviction ratio remains uniform across all attention heads, restricting head-level flexibility. Flexibility is crucial in eviction strategies because models process varying types of data across different contexts, requiring the ability to selectively retain important tokens while discarding others. Without this adaptability, the model may inefficiently store redundant information, increasing memory usage and slowing down inference times.
Absence of global statistics. Eviction strategies that rely on accumulative attention scores, such as H2O, base their calculations on local statistics, which often results in attention bias issues (Oren et al., 2024). This local approach may misjudge a token’s importance, especially when the context spans long sequences. While several methods (Chen et al., 2024; Jo & Shin, 2024) have proposed solutions to mitigate this bias, they are still fundamentally rooted in local statistics, addressing the issue only partially. A more effective solution would incorporate global statistics to ensure that the eviction process is based on a more comprehensive understanding of the context, thereby reducing biases and improving the accuracy of token retention.
Inefficiency. Efficiency is another challenge for traditional methods such as H2O and FastGen (Ge et al., 2023). These methods operates in a token-by-token manner for each decoding step, sequentially deciding which key-value pairs to evict from the cache. Additionally, H2O determines which tokens to evict after computing the attention scores, meaning some computation is wasted on tokens that are later discarded. This sequential approach hinders efficiency, especially when handling large batches of data or when processing long sequences.
Our Attention-Gate mechanism is designed to address the aforementioned limitations.
@@ 3.3 Attention-Gate
The Attention-Gate (AG) is a lightweight, trainable module designed to determine which tokens in the KV-Cache of each attention-head should be retained or discardedAG is positioned before the Multi-Head Attention (MHA) layer and operates by generating eviction flags for each token, guiding both the computation of attention scores and the management of the KV-Cache.
Input and output. AG takes the hidden states as input and generates binary flags as output, one for each token in every attention head. These flags indicate whether the corresponding token’s key-value pairs should be stored or discarded in the KV-Cache. Specifically, the input is a hidden state matrix $X\in\mathbb{R}^{n\times d}$, where $n$ represents the sequence length, and $d$ denotes the dimensionality of the hidden states. The output consists of independent binary decisions for the retention or eviction of each token’s key-value pairs across all attention heads in the MHA module.
Local or global? The structure of AG offers two choices: using either local or global information to guide the eviction policy. 1. Local information: The simplest approach leverages a linear layer where each token only considers its own hidden state without taking other tokens in the sequence into account. This method is computationally efficient and straightforward, making it appealing for its simplicity. 2. Global information: On the other hand, a more comprehensive approach adopts an attention-like structure, where tokens can aggregate information from across the entire sequence. This enables the eviction policy to reflect the global in-context information, making decisions that are informed by the broader context. Naturally, this method is computationally heavier than the local approach, but it offers a more accurate reflection of the sequence’s token importance.
In our design, we opted for the attention-like structure because global context is essential for making effective eviction decisions. As demonstrated in Section 4.4, using an attention-like structure significantly outperforms the linear layer, as the latter struggles to properly evict tokens. This outcome is expected, as determining whether a token is redundant or crucial for inference often requires a global understanding of the sequence. A local approach, like a linear layer focusing on individual tokens, lacks the broader perspective necessary to make accurate decisions about which tokens are important. This finding also underscores the limitations of methods like H2O, which rely on local statistics rather than global ones, as mentioned in Section 3.2.
Softmax or Sigmoid? After gathering information through the attention-like structure, the next step is to compute the eviction probability for each token. Common activation functions such as Softmax and Sigmoid can be used to generate these probabilities.
In our case, we choose the Sigmoid function for a key reason: while Softmax normalizes probabilities across the entire sequence, introducing competition between tokens, Sigmoid treats each token independently. This independence is crucial because we do not want a token’s likelihood of being evicted to depend on the eviction probabilities of other tokens in the sequence.
@@
Eviction flags. After computing the eviction probabilities for each token, we apply a threshold $\tau$. If the probability of retaining a token exceeds $\tau$, that token is retained in the KV-Cache; otherwise, it is discarded. In the attention matrix, the columns corresponding to evicted tokens are masked out. However, the diagonal elements, where a token attends to itself, are always preserved. This guarantees that each token continues to interact with its own key.

### v3

@@ Abstract
The KV-Cache technique has become the standard for the inference of large language models (LLMs). Yet, it is widely criticized that KV-Cache can become a bottleneck of the LLM inference system. This paper enables a novel dynamic KV-Cache eviction policy by injecting a lightweight module called Attention-Gate to the model. It accepts the global context as input and yields eviction flags for each token. The self-attention modules in the model proceed according to the flags and cache only a subset of the KV states for next token prediction. The Attention-Gates can yield various flags for different heads and layers and be easily tuned on top of a pre-trained LLM via continual pre-training or supervised fine-tuning. The computational and memory overhead introduced by Attention-Gates can be minimal. We empirically evaluate the proposed approach across multiple scenarios, showing that effective eviction of redundant tokens can not only improve efficiency but also enhance performance.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) (Dubey et al., 2024; Team et al., 2024; Chiang et al., 2023) have achieved remarkable success across a wide range of tasks. A key technique that has enabled efficient LLM inference is KV-Cache, which stores transient attention keys and values to avoid recomputation. However, as the size of LLMs continues to increase and the demand for handling long-context queries grows, the KV-Cache has emerged as a significant bottleneck. Storing attention states for numerous tokens can lead to considerable memory overhead and data transfer among the memory hierarchy results in substantially increased inference time.
@@
To overcome these challenges, we introduce a parameterized KV-Cache eviction mechanism named Attention-Gate (AG), to perform reliable in-context eviction. AGs are positioned before self-attention layers within the model. They take a sequence of token features as input and generate eviction flags for the tokens, indicating whether the token should be excluded from subsequent self-attention computations. Tokens that are evicted do not require their KV states to be cached. AGs can be seamlessly integrated into pre-trained LLMs and tuned by minimizing the language modeling loss. Ideally, AGs can automatically learn to discern the most relevant tokens for the current context without manual intervention. In practice, we can implement the AG as a self-attention layer with much fewer heads than the original model (e.g., 4 v.s. 32). This way, the parallel computational capabilities of the hardware can be harnessed to minimize the extra overhead introduced by AGs.
AG is empirically shown to enjoy high training efficiency, e.g., only four NVIDIA 4090 GPUs and a dataset of 5,000 samples are required for continual pre-training when applying AGs to LLaMA2-7B (Touvron et al., 2023). This alleviates concerns about the computational overhead related to trainable eviction strategies (Zhang et al., 2024; Chen et al., 2024) and amplifies the merits of our approach over existing training-free approaches. As illustrated in Figure 1, AG generates different eviction strategies across different layers and attention-heads for different tokens, demonstrating its adaptability to the diverse requirements of each component in the model.
To validate the effectiveness of our method, we conduct extensive experiments across multiple benchmarks. After efficient continual pre-training (CPT), our approach outperforms traditional training-free eviction strategies, such as StreamingLLM and H2O, in both accuracy and token eviction rates. In supervised fine-tuning (SFT), our method not only evicts a significant number of redundant tokens but also maintains or surpasses the performance of LoRA-finetuned LLMs. For example, on the RTE dataset (Bar-Haim et al., 2006), our approach improves accuracy by 13.9% while evicting 62.8% of tokens, demonstrating that selective token eviction can enhance performance. In summary, the Attention-Gate mechanism provides a scalable and efficient solution for KV-Cache management, addressing the limitations of traditional training-free methods.
@@ 2 Related Work
Traditional KV-Cache eviction strategies. To address both memory and computational challenges, KV-Cache eviction has emerged as an effective strategy. Existing approaches predominantly rely on parameter-free heuristics.
@@
Strategies using accumulative attention scores offer more flexibility by dynamically identifying important tokens. For instance, SpAtten (Wang et al., 2021) employs Accumulative Attention Scores (A2S), which sum the softmax outputs for each token to measure its importance. This approach allows selective token pruning in subsequent layers, effectively reducing computational complexity without the need for retraining. H2O (Zhang et al., 2024) extends this concept to decoder-based models, using local A2S statistics for adaptive eviction in autoregressive generation. However, H2O suffers from the attention bias issue (Oren et al., 2024), particularly in long-context inputs. Several follow-up works have aimed to address this limitation. NACL (Chen et al., 2024) introduces random eviction to mitigate attention bias, while A2SF (Jo & Shin, 2024) incorporates a Forgetting Factor. However, none of these approaches fully resolve the underlying problem. Despite these limitations, some studies (Adams et al., 2024) suggest that H2O remains an optimal solution in many scenarios.
@@ 3.1 Preliminary
Multi-Head Attention (MHA) (Vaswani et al., 2017) is a core component of the Transformer architecture, as used by most LLMs. MHA enables the model to capture dependencies across different tokens in a sequence. Specifically, for an input sequence $X\in\mathbb{R}^{n\times d}$, where $n$ represents the sequence length and $d$ denotes the dimensionality of the hidden states, the output of MHA is computed as:
$\text{MHA}(X)=\left[\text{H}_{1}(X),\text{H}_{2}(X),\dots,\text{H}_{h}(X)\right]W^{O}\ ,$ (1)
where $\{\text{H}_{i}\}_{i=1}^{h}$ refers to the $h$ attention heads and
$\text{H}_{i}(X)$ $=\text{Attn}\left(XW^{Q}_{i},XW^{K}_{i},XW^{V}_{i}\right)$
$=\text{Attn}\left(Q_{i},K_{i},V_{i}\right)$ (2)
$=\text{Softmax}\left(\frac{Q_{i}K_{i}^{\top}}{\sqrt{d_{k}}}-\text{INF}(\mathbbm{1}-M_{i})\right)V_{i}=A_{i}V_{i}.$
Here, $W^{Q}_{i},W^{K}_{i}\in\mathbb{R}^{d\times d_{k}}$, $W^{V}_{i}\in\mathbb{R}^{d\times d_{v}}$, and $W^{O}\in\mathbb{R}^{hd_{v}\times d}$ are learned projection matrices. INF is a large constant, $\mathbbm{1}$ is a matrix of ones, $M_{i}$ is the mask applied to head Hi, and $A_{i}$ represents the attention scores for head Hi.
KV-Cache is employed during the inference of auto-regressive transformers, which stores the key and value information from previous time steps, allowing efficient reuse and reducing recomputation. The inference process can be divided into two stages: prefilling and decoding.
In the prefilling stage, the input sequence $X^{(\leq n)}=\left[x^{(1)},x^{(2)},\dots,x^{(n)}\right]\in\mathbb{R}^{n\times d}$ passes through MHA, and the corresponding key-value pairs $K^{(\leq n)}_{i}$ and $V^{(\leq n)}_{i}$ for head Hi are stored in KV-Cache. These are expressed as:
$K_{i}^{(\leq n)}=\left[k_{i}^{(1)},\cdots,k_{i}^{(n)}\right],\ V_{i}^{(\leq n)}=\left[v_{i}^{(1)},\cdots,v_{i}^{(n)}\right]\ ,$
where $k_{i}^{(t)}=x^{(t)}W_{i}^{K}$ and $v_{i}^{(t)}=x^{(t)}W_{i}^{V}$. After prefilling, the next token $x^{(n+1)}$ is generated.
In the decoding stage, $x^{(n+1)}$ is input to generate $x^{(n+2)}$ for the first step. During this process, only $k_{i}^{(n+1)},v_{i}^{(n+1)}$ need to be computed for head Hi. These are then concatenated with the cached $K_{i}^{(\leq n)}$ and $V_{i}^{(\leq n)}$ to form $K_{i}^{(\leq n+1)}$ and $V_{i}^{(\leq n+1)}$, which are used to complete the current MHA computation and update the KV-Cache. The process repeats token by token until the sequence generation is complete.
KV-Cache plays a critical role in improving the efficiency of LLM inference. However, the size of the KV-Cache grows with the input sequence length, leading to substantial memory overhead. Efficiently managing KV-Cache while maintaining model performance has become a key challenge in scaling LLMs to longer contexts.
@@ 3.2 Limitations of Traditional Eviction Strategies
Lack of flexibility. Static KV-Cache eviction strategies, such as those used in StreamingLLM (Xiao et al., 2024), lack adaptability across key dimensions, including token-specific, attention-head-specific, layer-specific, task-specific, and model-specific contexts. H2O (Zhang et al., 2024) addresses some of these limitations by introducing token-level and head-level adaptability. However, it still applies a uniform eviction ratio across all attention heads, which limits its granularity at the head level. This isotropic approach overlooks the significant variation in attention scores across different heads, as demonstrated by FastGen (Ge et al., 2023). This observation underscores the need for eviction strategies that are adaptive across multiple dimensions to better handle the diverse contexts processed by the model. Without finer-grained flexibility, existing training-free strategies risk retaining unnecessary information, leading to higher memory usage and reduced efficiency.
Absence of global statistics. KV-Cache eviction strategies should ideally be context-aware, as the importance of the same token can vary significantly depending on the surrounding context. To accurately discard redundant or less important tokens, it is necessary to consider the global statistics of the context. Strategies relying on accumulative attention scores, such as H2O, NACL, and A2SF (Chen et al., 2024; Jo & Shin, 2024; Zhang et al., 2024), are primarily based on local statistics of the context. While methods like NACL and A2SF attempt to mitigate the limitations of local context by introducing various adjustments to reduce the misjudgment of token importance and the resulting biased retention (Oren et al., 2024), they fail to address the root issue. The core problem lies in the lack of consideration of the global context when determining which tokens to evict.
Inefficiency. Methods like H2O and FastGen (Ge et al., 2023) are inefficient due to their sequential, token-by-token eviction processes at each decoding step. H2O, for example, computes attention scores before deciding which tokens to evict, wasting computation on soon-to-be discarded tokens.
Our method tries to address the aforementioned limitations.
@@ 3.3 Attention-Gate
The Attention-Gate (AG) is a lightweight, trainable module positioned before the MHA layer, generating eviction flags for the tokens. The flags determine which tokens in the KV-Cache of each head of the MHA should join the computation of attention scores and be retained in the KV-Cache.
Main architecture. AG consists of (i) an attention-like structure and (ii) a gating mechanism. (i) A modified version of the standard MHA in Equation 1, denoted as $\text{MHA}^{\prime}$, is used to facilitate global information exchange across all tokens in the sequence. To distinguish it from the vanilla MHA, all symbols in $\text{MHA}^{\prime}$ are marked with a prime ($\prime$). Notably, $\text{MHA}^{\prime}$ has much fewer heads compared to MHA, i.e., $h^{\prime}<h$, and its projection matrix $W^{O^{\prime}}\in\mathbb{R}^{h^{\prime}d^{\prime}_{k}\times h}$ differs slightly from $W^{O}\in\mathbb{R}^{hd_{v}\times d}$. (ii) The gating mechanism, denoted as G, is then introduced to enable adaptive eviction policies for attention heads within the MHA layer.
Input & output. AG takes the hidden states as input and generates binary flags as output, one for each token in every attention head. Specifically, given an input sequence $X\in\mathbb{R}^{n\times d}$, the output of AG is computed as:
$\text{AG}(X)=\text{G}\left(\text{MHA}^{\prime}(X),\tau\right)\ \in\{0,1\}^{n\times h},$ (3)
where $\tau$ is a hyperparameter. The gate G outputs 1 or 0 based on the score $s$ of a token obtained from MHA:
$\text{G}(s,\tau)=\begin{cases}1,&\text{if }\text{Sigmoid}(s)>\tau\\ 0,&\text{otherwise}\end{cases}\ .$ (4)
The values $k_{i}^{(t)}$ and $v_{i}^{(t)}$ for token $x^{(t)}$ and attention-head Hi are retained if $\text{AG}(X)_{i}^{(t)}=1$; otherwise, they are discarded. Based on this, the mask $M_{i}=\big{[}m_{i}^{(j,t)}\big{]}$ for the attention scores $A_{i}$ of head Hi in Section 3.1 is defined as:
$m_{i}^{(j,t)}=\begin{cases}0,&\text{if }j<t\\ 1,&\text{if }j=t\\ \text{AG}(X)_{i}^{(t)}\ \ ,&\text{otherwise}\end{cases}\ .$ (5)
@@
In the attention matrix, the columns corresponding to evicted tokens are masked out. However, the diagonal elements, where a token attends to itself, are always preserved. This ensures that each token maintains interaction with its own key, guaranteeing self-attention remains intact.
In this way, the AG module selectively determines which tokens are retained or discarded for each attention head, based on the global information captured by $\text{MHA}^{\prime}$ and the gating mechanism.
3.4 Design Rationale
Local or global? Our AG design, introduced in Section 3.3, utilizes a multi-head attention mechanism (i.e. MHA) that inherently incorporates global information. This approach allows the model to aggregate context across the entire sequence, enabling eviction decisions to accurately reflect the broader in-context information. Such a design is particularly effective for assessing the importance of tokens, as determining redundancy or relevance often requires a global understanding of the sequence.
Alternatively, a simpler approach can be used, such as employing a linear transformation to generate eviction flags. This method relies solely on local information, where each token only considers its own hidden state without incorporating information from other tokens in the sequence. While the local approach has the advantage of being computationally cheaper, as shown in Table 3, its performance and efficiency are not guaranteed. This limitation highlights the challenges faced by methods rely on local statistics of the context, as discussed in Section 3.2.
Softmax or Sigmoid? In our design, as described in Equation 4, we use the Sigmoid function to calculate the eviction probability for each token. However, it is worth noting that Softmax is another common choice for generating probabilities, as seen in applications like the gating mechanisms of Mixture of Experts in LLMs.
While both functions are viable, they offer different trade-offs. The Sigmoid function, as used in our design, computes probabilities independently for each token. This independence is important because it ensures that a token’s eviction likelihood is not directly influenced by the probabilities of other tokens in the sequence, allowing for a more flexible and precise eviction policy.
On the other hand, Softmax normalizes the probabilities across the entire sequence, which introduces competition between tokens. This can be beneficial in scenarios where the relative importance of tokens needs to be considered. However, in the context of eviction, where we want to assess each token individually without direct dependencies on others, Softmax might not be the ideal choice.